Find the equations of tangent and normal to the curve y=4x-x^2 at (1,3)
1
Expert's answer
2020-06-08T18:24:51-0400

Equations of tangent line and normal line can be found using:

"y-y_0=y'(x_0) (x-x_0) -" tangent line

"y-y_0=-\frac{1}{y'(x_0)}(x-x_0) -" normal line

So, all we need is to find "y'(x_0)"


"x_0=1\\y_0=3"

"y'(x)=(4x-x^2) '=4-2x", then

"y'(x_0) =y'(1) =4-2*1=2"


Finally:

"y-3=2(x-1) \iff y=2x+1" — tangent line

"y-3=-\frac{1}{2}(x-1) \iff y=-\frac{1}{2}x+\frac{7}{2}" — normal line